OpenStudy (anonymous):

solve by completing the square.x squared -9x=-14. please show work :)

OpenStudy (anonymous):



OpenStudy (lgbasallote):

x^2 - 9x = -14 x^2 - 9x + (9/2)^2 = -14 + (9/2)^2 x^2 -9x + 81/4 = -14 + 81/4 (x-9/2)^2 = 81/4 - 56/4 (x-9/2)^2 = 25/4 x - 9/2 = sqrt (25/4) x - 9/2 = +/- 5/2 x = 9/2 +/- 5/2 x = 14/2; x = 4/2 x=7 x = 2

OpenStudy (dumbcow):

take half of 2nd term and square it (-9/2)^2 = 81/4 add this to both sides --> x^2 -9x + 81/4 = -14 +81/4 the left side can now be factored into perfect square --> (x-9/2)^2 = 25/4 take square root --> x-9/2 = +-5/2 --> x = 9/2 +- 5/2 --> x = 7 or 2